OpenStudy (rane):

Mr Smith's small mathmetics class had percentage exams marks totalling 280. If the class exam mean was 40% ,how many students were in his class ?

OpenStudy (nirmalnema):

7 students were there....i think.. total no of students be x so there total marks is 280 280/ no.of students = mean.. 280/x=40 x=7

OpenStudy (rane):

how did u get 7 ? is that your guess?

OpenStudy (nirmalnema):

no... 280=40x x=280/40 x=7

OpenStudy (rane):

so u just divided by the percentage ?

OpenStudy (nirmalnema):

yup....is it correct..??

OpenStudy (rane):

nop

OpenStudy (nirmalnema):

so what is the correct ans..??

OpenStudy (rane):

idk, but its definitely not that small

OpenStudy (nirmalnema):

ok..

OpenStudy (rane):

and also your working is wrong u divided 280 by 40 = 7 but its not just "40 ", its 280 by 40% = 700

OpenStudy (nirmalnema):

hmm..

OpenStudy (nirmalnema):

40% is exam marks mean..??

OpenStudy (rane):

yh, class exam mean was 40%

OpenStudy (anonymous):

'Mr Smith's small mathematics class' 'small mathematics class' 'small' Yes it is 7 people don't be so sceptical about maths answers, again you can have people eating 200 melons a day

OpenStudy (rane):

@LS1088 can u kindly show me how did u get "7" and if its the same working as @nirmalnema then its wrong

OpenStudy (rane):

u guys are saying that 280 divided by 40 = 7 BUT its not just 40 its 40%

OpenStudy (nirmalnema):

no its not wrong 40% marks out of 100 = 40 so my working is correct//

OpenStudy (anonymous):

it is the same working, by mean of 40% I am certain the mean test score is not 0.4 as the total test marks would be 100

OpenStudy (nirmalnema):

x/100 * 100 =40 so x = 40

OpenStudy (nirmalnema):

yup ..@LS1088 is correct. n me too

OpenStudy (rane):

well, i did out 7 as the answer and its says its wrong

OpenStudy (rane):

put*

OpenStudy (anonymous):

Then is the question here the complete one?

OpenStudy (rane):

yh it is

OpenStudy (nirmalnema):

it is the only case when u get different ans..when the total no of marks of the paper is not 100

OpenStudy (rane):

well, idk since it says percentage exam marks totalling was 280

OpenStudy (nirmalnema):

hmm....